# Some remarks on real algebraic maps which are topologically special generic maps and generalize the canonical projections of the unit spheres

### v5

@@ Abstract.
In Proposition 2, if the interior unit sphere bundle and the boundary linear bundle for ${\bar{f}}_{N}$ are trivial and the identification bundle isomorphism for ${\bar{f}}_{N}$ is given as the product map of the identity map (identification) between the base spaces and the identity map on the fiber $S^{m-1}=\partial D^{m-n+1}$, then it is called a product-organized special generic map.
[20] has first introduced such a class of special generic maps. We first name the class in this way in [17]. Of course in our present paper, we do not need to understand this notion or [17, 20] well. We do not assume their arguments or results.
However, we note that canonical projections of unit spheres are of such a class. We consider a manifold represented as a connected sum $\sharp(S^{k_{j}}\times S^{m-k_{j}})$ of finitely many manifolds $S^{k_{j}}\times S^{m-k_{j}}$ where $m\geq n\geq 2$ and $11\leq k_{j}\leq n-$ 1 are integers. We choose the connected sum in the smooth category. We can reconstruct a product-organized special generic map into ${\mathbb{R}}^{n}$ there from a natural embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ whose image is diffeomorphic to the boundary connected sum $\natural(D^{k_{j}}\times S^{n-k_{j}})$ of finitely many manifolds $D^{k_{j}}\times S^{n-k_{j}}$ taken in the smooth category. This gives another simplest example.
We add the following new result. This is also a kind of remarks on Main Theorem 1.
@@
STEP 1 Obtaining another nice compact manifold isotopic to the original embedding.
@@
By considering a small isotopy considered in the class $C^{1}$ starting from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$, we have another smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n},0}:\bar{N}\rightarrow{\mathbb{R}}^{n}$. We can also consider the isotopy in the class $C^{a}$ where $a$ is an arbitrary integer greater than $1$ according to [23]. For this approximation, see also [4, 8].
This fixes the manifold except a small collar neighborhood $N(\partial\bar{N})$ of the boundary $\partial\bar{N}$. Let ${\bar{N}}_{0}\subset{\mathbb{R}}^{n}$ denote the resulting image ${\bar{f}}_{{\mathbb{R}}^{n},0}(\bar{N})$. Let ${\partial}_{j}{\bar{N}}_{0}$ denote all connected components of the boundary $\partial{\bar{N}}_{0}$, indexed by all positive integers $j=1,\cdots,l\geq 1$ smaller than or equal to $l$. We may also regard that the manifold ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$ (by considering a suitable small isotopy).
In addition, we can also get the embedding and the subset with the following properties.
@@
By considering the partial derivative for some variable $x_{j_{0}}$ in $x=(x_{1},\cdots x_{n})$, we have a value which is not 00. We discuss this more precisely. If $x\in{\partial}_{i_{0}}{\bar{N}}_{0}$ holds, then the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ is not 00 for some variable $x_{j_{0}}$, since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary. Remember that this is also obtained via an isotopy in the $C^{1}$ class. For the variable $x_{j_{0}}$, the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ is the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$. This is thanks to elementary formula in calculus and the property of the zero sets of our polynomials. Remember that $f_{{\partial}_{j}}(x)$ is not 00 for any $j$ except $j=i_{0}$. The value of the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ is not 00.
@@
We can have a small collar neighborhood $N(\partial{\bar{N}}_{0})$ of the boundary $\partial{\bar{N}}_{0}$, regarded as $(\partial{\bar{N}}_{0})\times[0,1]\subset N(\partial{\bar{N}}_{0})$. More precisely, we can have one and consider our identification in the following way thanks to arguments related to ”Discussion 14” in [21].
@@
This means that around $N(\partial{\bar{N}}_{0})$, $f_{0}$ is a special generic map and represented as the product map of a Morse function for a natural height of an ($m-n+1$)-dimensional disk and the identity map on $\partial{\bar{N}}_{0}$ (for suitable local coordinates).
Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ in $M_{0}$, remember that locally $M_{0}$ is represented as the graph of a canonically obtained smooth function. Some $y_{j}$ is regarded as the variable of the target of the function here.
We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{1}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule in the case $y$ is not the origin. This is a diffeomorphism. We map each point of the form $(x,0)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. Thus we have a diffeomorphism from $({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ onto $({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$. By the restriction, we also have another diffeomorphism $(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ onto $({\rm Int}\ {\bar{N}}_{0})\times S^{m-n}\subset({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$.
These arguments imply that $f_{0}$ is a product-organized special generic map.
@@
By considering the isotopy (in the $C^{1}$ class) in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}:M_{0}\rightarrow{\mathbb{R}^{n}}$ enjoying all desired properties in Main Theorems 1 and 2.
@@
We show Main Theorem 3. This is a kind of generalizations of Main Theorem 2.
Before this, we review several known facts on topologies and differentiable structures of spheres and disks shortly. See [35] and see also [13, 47, 48] for example. A smooth manifold which is homeomorphic to $S^{k}$ ($k\geq 1$) is diffeomorphic to $S^{k}$ for $k=1,2,3,5,6$ for example. It is well-known that smooth manifolds homeomorphic to $S^{k}$ are classified via abstract and sophisticated theory for $k\geq 7$. If a smooth manifold which is homeomorphic to $S^{k}$ is smoothly embedded into ${\mathbb{R}}^{k+1}$ for $k\geq 1$ with $k\neq 4$, then this is diffeomorphic to $S^{k}$. A smooth manifold ”which is homeomorphic to $S^{4}$ and not diffeomorphic to $S^{4}$” is still undiscovered. A smooth manifold which is homeomorphic to $D^{k}$ and whose boundary is diffeomorphic to $\partial D^{k}=S^{k-1}$ is known to be diffeomorphic to $D^{k}$ unless $k=4$. A smooth manifold which is homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ unless $k=4,5$. Remember that ”${\Delta}^{k}$ which is homeomorphic to $D^{k}$ and is not diffeomorphic to $D^{k}$” is still undiscovered. If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$ or $5$.
@@
    $F_{S^{m-n}}(y,t)=F_{0,S^{m-n}}(t)-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some polynomials $F_{0,S^{m-n}}$ and $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ and $(y,t)\in{\mathbb{R}}^{m-n+1}\times\mathbb{R}$.
@@
    We have a diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ mapping each sphere ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a diffeomorphism where $a$ and ${\Sigma}^{m-n}$ are a suitably chosen positive real number, and a suitably chosen smooth manifold homeomorphic to $S^{m-n}$, respectively. The zero set ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ of the polynomial is non-singular and it does not contain the origin $0\in{\mathbb{R}}^{m-n+1}$. It is also homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
Then we can have a statement of Main Theorem 1 by defining
@@
for a sufficiently large $T>0$ and real polynomials $f_{{\partial}_{j}}$ in our proof of Main Theorems 1 and 2.
@@
    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example. We have a smooth bundle whose projection is the composition of the restriction of $q_{f_{0}}$ to the preimage of a suitable small collar neighborhood $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and the boundary of whose fiber is homeomorphic to $S^{m-n}$. Furthermore, the smooth bundle is a trivial smooth bundle.
@@
    The non-singular real algebraic manifold $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ homeomorphic to a manifold obtained in the following steps. Here ${\Delta}^{k}$ is for a $k$-dimensional smooth manifold homeomorphic to $D^{k}$ ($k\geq 1$). This is diffeomorphic to $D^{k}$ for $k\neq 4,5$. This is also diffeomorphic to $D^{k}$ for $k\neq 4$ if the boundary is diffeomorphic to $S^{k-1}$.
@@
We can argue as STEP 1 in the proof of Main Theorems 1 and 2. We can abuse the notation there and we do.
We choose a sufficiently large number $T>0$ such that the maximal value of $\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$ is smaller than the given positive number $a>0$. We choose another positive number $T_{0}<a$. We define $M_{0}$ as presented in the statement of Main Theorem 3.
@@
We present arguments similar to ones of STEP 2 in the proof of Main Theorems 1 and 2. Here we also respect the assumptions (1)–(4) here.
@@
by our definitions similarly. We show that $M_{0}$ is also a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
@@
Around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function. We discuss this more precisely. We consider the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some variable $y_{j}$. According to the property (4), ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$. It is assumed in (2) that $F_{0,S^{m-n}}(t)>0$ for $t>0$. We can know that the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ is not 00 for some variable $y_{j}$.
@@
Around a point of the form $(x,y)\in((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, we have a partial derivative as in STEP 2 in the proof of Main Theorems 1 and 2. The value is the product of the derivative of $F_{0,S^{m-n}}$ at 00, which is not 00 from the assumption (2), $\frac{1}{T}$, and the value in the case of Main Theorems 1 and 2. The value in the case of Main Theorems 1 and 2 means the value of the partial derivative of ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))$ for the variable $x_{j_{0}}$. The resulting value is not 00.
@@
We can argue similarly to see that $M_{0}$ is a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
@@
We present an argument similar to one in STEP 3 in the proof of Main Theorems 1 and 2. For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$. For example, we can consider this as the collar neighborhood of STEP 3 in the proof. We regard this as $(\partial{\bar{N}}_{0})\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $(\partial{\bar{N}}_{0})\times\{0\}$ are identified by the correspondence identifying $x$ with $(x,0)$.
@@
We construct a homeomorphism from the manifold $M_{0}$ onto another smooth manifold $M$.
@@
First we map $(\partial{\bar{N}}_{0})\times\{0\}$ to $(\partial{\bar{N}}_{0})\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$. We use the natural identification mapping $(x,0)$ to $(x,0)$ here. According to the property (3) for example, we have $((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}=(\partial{\bar{N}}_{0})\times\{0\}$.
@@
Hereafter, let ${\rm Pr}_{{\Sigma}^{m-n},a,1}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\Sigma}^{m-n}$ denote the projection to the first component and ${\rm Pr}_{{\Sigma}^{m-n},a,2}:{\Sigma}^{m-n}\times(0,a]\rightarrow(0,a]$ the projection to the second component.
@@
We can map each point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ to
$(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),T_{0}))$

### v6

@@ Abstract.
In Proposition 2, if the interior unit sphere bundle for ${\bar{f}}_{N}$ and the boundary linear bundle for ${\bar{f}}_{N}$ are trivial and the identification bundle isomorphism for ${\bar{f}}_{N}$ is the product map of the identity map on the base space (or the identification between the base spaces) and the identity map on the fiber $S^{m-1}=\partial D^{m-n+1}$, then it is called a product-organized special generic map.
[20] has first introduced such a class of special generic maps. We first name the class in this way in [17]. Of course in our present paper, we do not need to understand precise arguments or results from [17, 20] well.
We note the canonical projection of the unit sphere is of such a class. We consider a manifold represented as a connected sum $\sharp(S^{k_{j}}\times S^{m-k_{j}})$ of finitely many manifolds $S^{k_{j}}\times S^{m-k_{j}}$ where $m\geq n\geq 2$ and $11\leq k_{j}\leq n-$ 1 are integers. We take the connected sum in the smooth category. We can reconstruct a product-organized special generic map into ${\mathbb{R}}^{n}$ there from a natural embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ whose image is diffeomorphic to the boundary connected sum $\natural(D^{k_{j}}\times S^{n-k_{j}})$ of finitely many manifolds $D^{k_{j}}\times S^{n-k_{j}}$ taken in the smooth category. This gives another simplest example.
We have the following by the definition and the structures of the maps immediately. In the following, we abuse the notation of Propositions 1 and 2 or its natural variant.
Proposition 3.
Let $f_{1}:M\rightarrow N$ be a product-organized special generic map and for the induced smooth immersion $\bar{f}:=\bar{f_{1}}:W_{f_{1}}\rightarrow N$ and another smooth immersion ${{\bar{f}}_{N}}^{\prime}:W_{f_{1}}\rightarrow N$, let there exist a homeomorphism ${\phi}_{N_{1,2}}:N\rightarrow N$ satisfying the relation ${{\bar{f}}_{N}}^{\prime}={\phi}_{N_{1,2}}\circ\bar{f_{1}}$. In this situation, there exists a product-organized special generic map $f_{2}:M\rightarrow N$ satisfying the relation $f_{2}={\phi}_{N_{1,2}}\circ f_{1}$ and $\bar{f_{2}}={{\bar{f}}_{N}}^{\prime}$.
In short, Remark 1 (2) gives Proposition 3. We explain some, more precisely. First, from $f_{1}$, we have two kinds of bundles as in Proposition 1. From these two bundles and the identification between the boundaries, we obtain the three objects for reconstruction of a special generic map in Proposition 2 canonically. We can consider a smooth surjection as in Remark 1 (2), and after that, we compose the immersion ${{\bar{f}}_{N}}^{\prime}$ to have our desired special generic map $f_{2}$.
The following new result is a kind of our remark on Main Theorem 1.
@@
STEP 1 Obtaining another nice compact manifold which is isotopic to the original embedding and whose boundary consists of connected components each of which is the zero set of some real polynomial function and non-singular.
@@
By considering a suitable small isotopy of the class $C^{1}$ starting from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$, we have another nice smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n},0}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ in the following. Before explaining the new embedding, we give some small remark. We can also consider the isotopy in the class $C^{a}$ where $a$ is an arbitrary integer greater than $1$ according to [23]. For this approximation, see also [4, 8]. This can be also chosen as an isotopy fixing the manifold except a small collar neighborhood $N(\partial\bar{N})$ of the boundary $\partial\bar{N}$.
Let ${\bar{N}}_{0}\subset{\mathbb{R}}^{n}$ denote the resulting image ${\bar{f}}_{{\mathbb{R}}^{n},0}(\bar{N})$. Let ${\partial}_{j}{\bar{N}}_{0}$ denote all connected components of the boundary $\partial{\bar{N}}_{0}$, indexed by all positive integers $j=1,\cdots,l\geq 1$ smaller than or equal to $l$. The manifold ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$. In addition, we can also get the embedding and the subset with the following properties.
@@
By considering the partial derivative for some variable $x_{j_{0}}$ in $x=(x_{1},\cdots x_{n})$, we have a value which is not 00. We discuss this more precisely. If $x\in{\partial}_{i_{0}}{\bar{N}}_{0}$ holds, then the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ is not 00 for some variable $x_{j_{0}}$, since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary. Remember that this is also obtained via an isotopy in the $C^{1}$ class. For the variable $x_{j_{0}}$, the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ is the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$. This is thanks to elementary formula in calculus and the property of the zero sets of our polynomial functions. Remember that $f_{{\partial}_{j}}(x)$ is not 00 for any $j$ except $j=i_{0}$. The value of the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ (for $x_{j_{0}}$) is not 00.
@@
We can have a small collar neighborhood $N(\partial{\bar{N}}_{0})$ of the boundary $\partial{\bar{N}}_{0}$, regarded as $(\partial{\bar{N}}_{0})\times[0,1]\subset N(\partial{\bar{N}}_{0})$. More precisely, we can have one and consider our identification in the following way thanks to arguments related to [21, Discussion 14].
@@
This means that around $N(\partial{\bar{N}}_{0})$, $f_{0}$ is a special generic map and represented as the product map of the Morse function for the natural height of the ($m-n+1$)-dimensional unit disk $D^{m-n+1}$ and the identity map on $\partial{\bar{N}}_{0}$ (for suitable local coordinates).
Around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, remember that locally $M_{0}$ is represented as the graph of a canonically obtained smooth function. Some $y_{j}$ is regarded as the variable of the target of the function here.
We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times({\mathbb{R}}^{m-n+1}-\{0\})$ to $(x,\frac{1}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$. This gives a diffeomorphism on $({\rm Int}\ {\bar{N}}_{0})\times({\mathbb{R}}^{m-n+1}-\{0\})$. We map each point of the form $(x,0)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. Thus we have a diffeomorphism from $({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ onto $({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$. By the restriction, we also have another diffeomorphism $(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ onto $({\rm Int}\ {\bar{N}}_{0})\times S^{m-n}\subset({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$.
The structures of the manifolds and maps and these arguments imply that $f_{0}$ is a product-organized special generic map.
@@
We check the existence of desired maps for Main Theorems 1 and 2. If we do not need to deform the embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ in STEP 1 and we can regard ${\bar{f}}_{{\mathbb{R}}^{n}}={\bar{f}}_{{\mathbb{R}}^{n},0}$, then we can have desired maps by putting the special generic map $f:=f_{0}$ and the homeomorphism $\phi$ as the identity map.
In general, we need to deform the embedding in STEP 1 and in the general case, our desired map $f$ is reconstructed by Propositions 2 and 3 and Definition 1 similarly from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$. For the original embedding, each connected component of the boundary of the image ${\bar{f}}_{{\mathbb{R}}^{n}}(\bar{N})$ may not be real algebraic (as a result $f$ may not be real algebraic). We can also know the existence of the homeomorphism $\phi$ for Main Theorems 1 and 2 by the structures of the manifolds and maps and Proposition 3, easily.
@@
We show a kind of generalizations for Main Theorems 1 and 2. Before this, we review several known facts on topologies and differentiable structures of spheres and disks shortly. See [35] and see also [13, 47, 48], for example. A smooth manifold which is homeomorphic to $S^{k}$ ($k\geq 1$) is diffeomorphic to $S^{k}$ for $k=1,2,3,5,6$ for example. It is well-known that smooth manifolds homeomorphic to $S^{k}$ are classified via abstract and sophisticated theory for $k\geq 7$. If a smooth manifold which is homeomorphic to $S^{k}$ is smoothly embedded into ${\mathbb{R}}^{k+1}$ for $k\geq 1$ with $k\neq 4$, then this is diffeomorphic to $S^{k}$. A smooth manifold ”which is homeomorphic to $S^{4}$ and not diffeomorphic to $S^{4}$” is still undiscovered. A smooth manifold which is homeomorphic to $D^{k}$ and whose boundary is diffeomorphic to $\partial D^{k}=S^{k-1}$ is known to be diffeomorphic to $D^{k}$ unless $k=4$. A smooth manifold which is homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ unless $k=4,5$. Remember that ”${\Delta}^{k}$ which is homeomorphic to $D^{k}$ and is not diffeomorphic to $D^{k}$” is still undiscovered. If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$ or $5$.
@@
    $F_{S^{m-n}}(y,t)=F_{0,S^{m-n}}(t)-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some real polynomials $F_{0,S^{m-n}}$ and $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ and $(y,t)\in{\mathbb{R}}^{m-n+1}\times\mathbb{R}$.
@@
    A positive number $a>0$ and a smooth manifold ${\Sigma}^{m-n}$ with a homeomorphism ${\phi}_{\Sigma}:{\Sigma}^{m-n}\rightarrow S^{m-n}$ are given. There exists a homeomorphism ${\Phi}_{F_{{S}^{m-n}},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ mapping each sphere ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a homeomorphism. Furthermore, the zero set ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ of the real polynomial function is non-singular for $t\in(0,a]$.
In this situation, by defining
@@
for a sufficiently large $T>0$ and real polynomials $f_{{\partial}_{j}}$ in our proof of Main Theorems 1 and 2, we have an $m$-dimensional smooth manifold homeomorphic to $M_{0}$ and the following maps, similar to those in Main Theorem 1, as follows.
  1. (5)
    A special generic map $f:M\rightarrow N:={\mathbb{R}}^{n}$ in Proposition 1 with $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}:W_{f}=\bar{N}\rightarrow{\mathbb{R}}^{n}$.
  2. (6)
    A real algebraic map $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ satisfying the relation $f_{0}\circ\Phi=\phi\circ f$ for some pair $(\Phi:M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms with the relations $\Phi(S(f))=S(f_{0})$ and $\phi(f(S(f)))=f_{0}(S(f_{0}))$.
@@
    We can abuse the notation in Proposition 1 replacing ”$f$” by ”$f_{0}$” for example. We have a smooth bundle whose projection is the composition of the restriction of $q_{f_{0}}$ to the preimage of a suitable small collar neighborhood $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$, and the boundary of whose fiber is homeomorphic to $S^{m-n}$. Furthermore, the smooth bundle is a trivial smooth bundle.
@@
    The non-singular real algebraic manifold $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ homeomorphic to a manifold obtained in the following steps. Here ${\Delta}^{k}$ is for a $k$-dimensional smooth manifold homeomorphic to $D^{k}$ ($k\geq 1$). This is diffeomorphic to $D^{k}$ for $k\neq 4,5$. This is also diffeomorphic to $D^{k}$ for $k\neq 4$ if its boundary is diffeomorphic to $S^{k-1}$.
@@
We can argue as ”STEP 1 in the proof of Main Theorems 1 and 2”. We can abuse the notation there and we do.
We choose a sufficiently large number $T>0$ such that the maximal value of $\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$ is smaller than the given positive number $a>0$. We define $M_{0}$ as presented in the statement of Main Theorem 3.
@@
We present arguments similar to ones of ”STEP 2 in the proof of Main Theorems 1 and 2”. Here we also respect the assumptions (1)–(4) here.
@@
by our definitions and the assumptions (1)–(3) similarly. The corresponding argument ”in the proof of Main Theorems 1 and 2” is naturally generalized.
@@
We check that $M_{0}$ is also a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
@@
Around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function. We discuss this more precisely. We consider the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some variable $y_{j}$. According to the property (4), ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular for $t\in(0,a]$. We have checked that the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ is not 00 for some variable $y_{j}$.
@@
Around a point of the form $(x,y)\in((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, we have a partial derivative as in ”STEP 2 in the proof of Main Theorems 1 and 2”. The value is the product of the derivative of $F_{0,S^{m-n}}$ at 00, which is not 00 from the assumption (2), $\frac{1}{T}$, and the ”value in the case of Main Theorems 1 and 2”. The ”value in the case of Main Theorems 1 and 2” means the value of the partial derivative of ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))$ for the ”variable $x_{j_{0}}$ there”. The resulting value is not 00.
@@
We have checked similarly that $M_{0}$ is a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
@@
We present an argument similar to one in ”STEP 3 in the proof of Main Theorems 1 and 2”. For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$. For example, we can consider this as the collar neighborhood of ”STEP 3 in the proof of Main Theorems 1 and 2”. We regard this as $(\partial{\bar{N}}_{0})\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $(\partial{\bar{N}}_{0})\times\{0\}$ are identified by the correspondence identifying $x$ with $(x,0)$.
@@
We construct a homeomorphism ${\Phi}^{\prime}$ from the manifold $M_{0}$ onto another new smooth manifold $M$.
@@
First we map $(\partial{\bar{N}}_{0})\times\{0\}$ to $(\partial{\bar{N}}_{0})\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$. We use the natural identification mapping $(x,0)$ to $(x,0)$ here. According to the assumption (3), we have $((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}=(\partial{\bar{N}}_{0})\times\{0\}$.
@@
Hereafter, let ${\rm Pr}_{{\Sigma}^{m-n},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\Sigma}^{m-n}$ denote the projection to the first component.
@@
Remember that our original homeomorphism ${\Phi}_{F_{S^{m-n}},a}$ is presented in the assumption (4): this maps ${\Sigma}^{m-n}\times(0,a]$ onto ${\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ mapping each homotopy sphere ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a homeomorphism. We can map each point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ to
$(x,g(t){\phi}_{\Sigma}\circ{\rm Pr}_{{\Sigma}^{m-n},a}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y))$